OpenStudy (anonymous):

Find the measure of each exterior angle of a regular pentagon. A. 36° B. 72° C. 108° D. 120°

jimthompson5910 (jim_thompson5910):

E = 360/n in this case, n = 5 since there are 5 sides E = exterior angle n = # of sides

OpenStudy (anonymous):

72 degrees @jim_thompson5910

OpenStudy (anonymous):

good.

OpenStudy (anonymous):

you got it now, just memorize the E=360/n formula and youre good to go. :)